Standart Normal Probabilities

Let Z be a standard normal random variable. Calculate the following probabilities using the calculator provided. Round your responses to at least three decimal places. P(Z >-2.11) P(Z 1.82) = P (-048<Z < 205) Clear Undo Help Next>> Explain
Answer #1

P(Z>-2.11) = 1- P(Z<= -2.11) = 0.983

P(Z<=1.82) = 0.965

P(-0.48<Z<2.05) = P(Z<2.05) - P(Z< -0.48) = 0.664